OpenStudy (anonymous):

Differentiate the following functions with respect to x: y = ln7x

OpenStudy (anonymous):

dy(u(x))/dx = (dy/du)(du/d) chain rule u=7x, y(u(x)) = ln(7x) dy/du = 1/7x, du/dx=7 dy/dx = 1/x or y = ln(7) + ln(x) derivative of constant is 0 d ln(x)/dx = 1/x